# First variation of flat traces on negatively curved surfaces

### v1

@@ 1. Introduction
The rigidity statement in Theorem 1.2 is weaker than marked length spectrum rigidity, since it assumes a smooth deformation. The point of the present paper is the distributional first variation formula for the flat trace, in particular the transport coefficient in Proposition 2.2. This coefficient does not seem to appear in the literature in this explicit form, and it is the mechanism that converts constancy of a zeta type invariant into infinitesimal constraints on the marked lengths. The appendices collect auxiliary calculations and are included for completeness.
@@ 2.1. Kernel and flat trace as a pullback and pushforward
2.2. Normal form near a closed orbit
@@ 5.1. Guillemin-Kazhdan variational identity
where $ds$ denotes the Hamiltonian time parameter along $y_{0}$.
@@
Let $\alpha$ be the canonical $1$-form on $T^{*}X$ and $\omega=d\alpha$ the canonical symplectic form. Let $H_{t}$ be the Hamiltonian vector field of $p_{t}$, characterized by
@@
Since $p_{t}$ is homogeneous of degree $1$ in $\xi$, Euler’s identity gives $\alpha(H_{t})=p_{t}$, and therefore
@@ 6.1. Explicit computation of the lower order terms in (7)
Let $\chi\in C_{c}^{\infty}(U)$ be a cutoff which equals $1$ in a smaller neighborhood of $\gamma$ and set $\chi^{\Psi}:=\chi\circ\Psi$. The localized flat trace is the distribution
@@
Negative curvature implies $\gamma$ is hyperbolic, hence $1$ is not an eigenvalue of $P_{\gamma}^{\,m}$ and $\det(I-P_{\gamma}^{\,m})\neq 0$. By the inverse function theorem, $F_{mL^{\#}_{\gamma}}$ is a diffeomorphism from a neighborhood of 0 onto a neighborhood of 0, and the distribution $\delta(F_{mL^{\#}_{\gamma}}(y))$ is computed by a change of variables: for any $\psi\in C_{c}^{\infty}(\mathbb{R}^{2})$ supported sufficiently close to 0,
@@ 6.2. The $SO(2)$ Fourier calculus on $S^{*}_{g}X$ and the $\pm 1$ mode reduction
Let $(X,g)$ be a closed oriented surface with Gaussian curvature $K<0$, and identify $S^{*}_{g}X\simeq SX$ via the musical isomorphism. Denote by $\pi:SX\to X$ the bundle projection. We use the standard canonical coframing on $SX$: there exist smooth $1$-forms $(\alpha,\beta,\psi)$ and smooth vector fields $(X,X_{\perp},V)$ on $SX$ uniquely specified by
@@
Commutators. For any $1$-form $\omega$ and vector fields $Y,Z$,
@@
In our application $f=\dot{p}|_{SX}=-\tfrac{1}{2}h(v,v)$, where $h=\dot{g}$ is a smooth symmetric $2$-tensor on $X$ and $v\in S_{x}X$ is the unit direction. To determine the $SO(2)$-Fourier support of $\dot{p}$, fix a local oriented $g$-orthonormal frame $(e_{1},e_{2})$ on $X$ and write the fiber angle coordinate $\theta$ so that

### v2

@@ 1. Introduction
The rigidity statement in Theorem 1.2 is weaker than marked length spectrum rigidity, since it assumes a smooth deformation. The point of the present paper is the distributional first variation formula for the flat trace, in particular the transport coefficient in Proposition $2.$ This coefficient does not seem to appear in the literature in this explicit form, and it is the mechanism that converts constancy of a zeta type invariant into infinitesimal constraints on the marked lengths. The appendices collect auxiliary calculations and are included for completeness.
@@ 2.1. Kernel and flat trace as a pullback and pushforward
$2.$ Normal form near a closed orbit
@@ 5.1. Guillemin-Kazhdan variational identity
where dsds denotes the Hamiltonian time parameter along $y_{0}$.
@@
Let $\alpha$ be the canonical 11-form on $T^{*}X$ and $\omega=d\alpha$ the canonical symplectic form. Let $H_{t}$ be the Hamiltonian vector field of $p_{t}$, characterized by
@@
Since $p_{t}$ is homogeneous of degree 11 in $\xi$, Euler’s identity gives $\alpha(H_{t})=p_{t}$, and therefore
@@ 6.1. Explicit computation of the lower order terms in (7)
Let $\chi\in C_{c}^{\infty}(U)$ be a cutoff which equals 11 in a smaller neighborhood of $\gamma$ and set $\chi^{\Psi}:=\chi\circ\Psi$. The localized flat trace is the distribution
@@
Negative curvature implies $\gamma$ is hyperbolic, hence 11 is not an eigenvalue of $P_{\gamma}^{\,m}$ and $\det(I-P_{\gamma}^{\,m})\neq 0$. By the inverse function theorem, $F_{mL^{\#}_{\gamma}}$ is a diffeomorphism from a neighborhood of 0 onto a neighborhood of 0, and the distribution $\delta(F_{mL^{\#}_{\gamma}}(y))$ is computed by a change of variables: for any $\psi\in C_{c}^{\infty}(\mathbb{R}^{2})$ supported sufficiently close to 0,
@@ 6.2. The $SO(2)$ Fourier calculus on $S^{*}_{g}X$ and the $\pm 1$ mode reduction
Let $(X,g)$ be a closed oriented surface with Gaussian curvature $K<0$, and identify $S^{*}_{g}X\simeq SX$ via the musical isomorphism. Denote by $\pi:SX\to X$ the bundle projection. We use the standard canonical coframing on SXSX: there exist smooth 11-forms $(\alpha,\beta,\psi)$ and smooth vector fields $(X,X_{\perp},V)$ on SXSX uniquely specified by
@@
Commutators. For any 11-form $\omega$ and vector fields $Y,Z$,
@@
In our application $f=\dot{p}|_{SX}=-\tfrac{1}{2}h(v,v)$, where $h=\dot{g}$ is a smooth symmetric 22-tensor on $X$ and $v\in S_{x}X$ is the unit direction. To determine the $SO(2)$-Fourier support of $\dot{p}$, fix a local oriented $g$-orthonormal frame $(e_{1},e_{2})$ on $X$ and write the fiber angle coordinate $\theta$ so that